Consider the total cost function TC= 120y-y^2+0.02y^2.

a. What level of y yields minimum cost per unit of y?

b. What is the relationship between MC and AVc at the above level of output?


1
Expert's answer
2021-07-19T13:44:54-0400

a.

"TC= 120y-y^2+0.02y^2\\
0=120-2y+0.04y\\
0=120-1.96y\\
y=61.22"

b.

The decreasing average variable cost (AVC) responds to growing factor returns, whereas the rising marginal cost (MC) responds to rising marginal product. As a result, the rising MC and dropping AVC indicate that the MC will be lower than the AVC.